# Localized Deviations from the CO–PAH Relation in PHANGS–JWST Galaxies: Faint PAH Emission or Elevated CO Emissivity?

### v1

@@ 4.1 Suppressed PAH Emission Scenario
As described in Section 2.1, the JWST Cycle 1 data provide coverage across multiple filters, enabling measurements of key PAH emission features (at 3.3, 7.7, and 11.3m) as well as the dust continuum (at 10m and 21m). Here, we investigate whether the outlier regions identified in the CO-to-7.7m PAH ratio persist when other PAH or dust continuum dominated bands are used. We also compare PAH properties between the abnormal and null-test regions.
4.1.1 CO-to-mid-IR ratio using other PAH and dust continuum emission
@@
Figure 2 shows the CO flux relative to that expected from other PAH- and dust continuum–dominated bands in NGC 4303, which is one of the galaxies with the most prominent CO-to-7.7m PAH anomalies. We estimate CO fluxes from 3.3 and 11.3m bands using the coefficients from R. Chown et al. (2025) ( and ), and from 10 and 21m bands using coefficients derived in this work following the same method ( and ).
@@
The abnormal regions identified in CO-to-7.7m PAH ratio (red) consistently show elevated ratios across all PAH and continuum bands, with the largest excess observed at 3.3m and the smallest at 10m. Other galaxies hosting abnormal regions also show a similar behavior. The significant enhancement across multiple bands indicates that 7.7m is not particularly suppressed and the anomaly is more likely due to overluminous CO emission rather than a physical mechanism that suppresses multiple PAH bands and dust continuum emission at the same time. It is also interesting to note that shows the closest agreement with , even though m traces mostly PAH emission while m traces mostly dust continuum. Further supporting this picture, Appendix A shows three–color HST composites (F814W, F438W, F555W; J. C. Lee et al., 2022) in which the outlier regions display dust extinction similar to their surroundings (see also H. Faustino Vieira et al. 2025). This suggests that abnormal regions exhibit comparable dust content compared to their surroundings, making simultaneous suppression of PAH and dust emission an unlikely explanation for the elevated CO/PAH ratios.
4.1.2 Difference in PAH properties
@@
PAH intensity ratios are sensitivity to molecule properties (e.g., size, asymmetry, composition, and charge state) and to local radiation field (S. Hony et al., 2001; F. Galliano et al., 2008; T. S. Y. Lai et al., 2020; A. Maragkoudakis et al., 2020). In particular, 11.3m/3.3m becomes lower as the spectral energy distribution of the illumination radiation becomes harder and PAH size distributions become smaller. The ratio of 11.3m/7.7m also decreases with harder radiation field and is very sensitive to changes of PAH charge variations (T. S. Y. Lai et al., 2022; J. Chastenet et al., 2023; B. T. Draine et al., 2021; B. S. Hensley & B. T. Draine, 2023b; D. Baron et al., 2024, 2025). Furthermore, the ratio of emission in PAH bands to the 21 m dust continuum is considered as a strong tracer of the fraction of dust mass in the form of PAH (J. Chastenet et al., 2019; J. Sutter et al., 2024; O. V. Egorov et al., 2023).
@@
In Figure 3, we show median and 1 distribution of PAH band ratios to assess whether PAH properties differ between our abnormal and null-test regions across Cycle 1 galaxies. The fraction of dust mass in PAHs, , estimated from , shows no significant difference between abnormal and null-test regions. We note that the abnormal and null test regions also do not overlap with massive star-forming sites, where is typically suppressed due to PAH destruction by stellar feedback (J. Chastenet et al., 2023; O. V. Egorov et al., 2023, 2025; J. Sutter et al., 2024).
@@
In the middle panel, we compare . While most galaxies follow a one-to-one relation, abnormal regions in NGC 1433 and NGC 4535 show somewhat lower compared to their null-test region, suggesting fragmentation into smaller grains or harder radiation field. In contrast, NGC 3627, shows a higher ratio in its abnormal region (see also (D. A. Dale et al., 2025)), possibly indicating PAH growth or coagulation onto dust grains, making 7.7m and 11.3m bands fainter (B. T. Draine et al., 2021; B. S. Hensley & B. T. Draine, 2023b). As most of these outlier regions lie along bar lanes with no signs of massive star formation, the differences in suggest that PAH size distributions may be altered by shocks or coagulation as gas and dust flow inward.
@@
Lastly, we compare PAH ratios. The , often used as a tracer of PAH ionization (e.g., J. D. T. Smith et al., 2007; T. S. Y. Lai et al., 2022), has also been shown to correlate with (D. Baron et al., 2024). This implies that the majority of the variation in ratios is driven by the changing shape of the illuminating spectral energy distribution, where the ratio becomes higher as the radiation from old stellar population dominates (high and high ) compared to the radiation from young massive stars (low and low ). In NGC 1433 and NGC 4321, we find somewhat higher in abnormal regions ( dex), which may indicate an older stellar population dominating the radiation and/or more neutral PAHs, and therefore the lack of PAH emission compared to CO. However, in most galaxies, abnormal and null-test regions agree within 1.
@@
While some abnormal regions show hints of altered PAH sizes or hardness of the radiation field, these trends are not consistent across systems. As the PAH ratios explored here are degenerate with multiple physical effects, additional spectroscopic tracers (e.g., H2 rotational and rovibrational lines, [Fe II] for shocks, and a wider suite of PAH features) will be essential to further investigate whether PAH emission is suppressed in these regions.
4.2 Elevated CO emissivity scenario
@@
Almost all abnormal regions lie along the dust lanes in bars, sites where inflowing gas experiences streaming motions and strong shocks due to intersecting orbital families, causing abrupt changes in velocity (W. W. Roberts et al., 1979; E. Athanassoula, 1992; M. C. Sormani et al., 2015; J. A. Sellwood & A. Wilkinson, 1993). These dynamical processes can elevate the CO velocity dispersion and increase the integrated intensity either through line broadening or by reducing the optical depth, which boosts the CO emissivity (A. D. Bolatto et al., 2013; M. Gong et al., 2020; Y.-H. Teng et al., 2023). In addition, shocks and strong shear can raise the excitation temperature and collision rates, populating higher- levels and further enhancing the CO flux for a given molecular gas mass.
@@
In this Section, to test if high CO-to-PAH ratio is due to elevated CO emissivity or excitation by comparing CO velocity dispersions and CO line ratios between outlier and null-test regions. If the high CO/PAH ratios are associated with higher velocity dispersions or elevated excitation temperature, it would indicate that the -to-PAH ratio between abnormal and null-test regions should remain similar and the CO–to– conversion factor () should be at least 10 times lower in abnormal regions compared to the regions with .
4.2.1 CO velocity dispersion
@@
Figure 4 shows the CO luminosity-weighted average velocity dispersion () as a function of the median CO-to-PAH ratio () of each abnormal region and null-test region, across the 20 galaxies hosting CO/PAH outliers. On average, the abnormal regions show two times higher and times higher compared to those in null-test regions555When all pixels are considered and not just the null-test region, the distribution of strongly peaks at 1.. The bottom panel shows as a function of the median CO-to-PAH ratio (). We obtain using the Y.-H. Teng et al. (2024) prescription,
with measured within each region. We also compare the distributions of and , where . As abnormal regions are associated with higher velocity dispersions and thus lower , the large offset observed between the abnormal and null-test regions in distribution decreases when is used. We also show how high (or how low the ) should be in order to explain these outliers solely by enhanced CO-emissivity, under the assumption of constant -to-PAH ratio (dotted line) between abnormal and null-test regions. While the abnormal regions are associated with higher velocity dispersion, it appears it should be even higher (three times more) to explain these outliers from enhanced emissivity.
The bottom panel of Figure 4 shows CO velocity dispersion () as a function of molecular gas surface density () for three galaxies where the enhancement in CO velocity dispersion is prominent. We also divide the data into logarithmically spaced bins in and then compute median and 1 scatter of in each bin for pixels in abnormal null-test regions. Within the range of where the abnormal and null-test regions overlap, the abnormal regions generally exhibit higher , indicating higher CO emissivity that can lead to abnormally high CO-to-PAH ratios.
In Appendix B, the same figure for all 20 galaxies hosting outlier regions is shown. While in some galaxies the velocity dispersions at a given agree within , almost half of them show differences exceeding this level, particularly NGC 1566, NGC 3351, NGC 4321, NGC 2566, NGC 2997, NGC 4569, NGC 4579, NGC 6300. We also note that when prescription from (Y.-H. Teng et al., 2024) is used, between abnormal and null-test regions overlaps much more significantly than when a constant Galactic is adopted (triangles). This again supports the view where abnormally high CO-to-PAH ratios are driven by elevated CO emissivity. The abnormal and null-test regions contain comparable amounts of PAH emission for a given amount of .
4.2.2 CO excitation from multi-transition lines and CO isotopologues
@@
Using CO isotopologues observations in their low rotational transitions, Y.-H. Teng et al. (2022) and Y.-H. Teng et al. (2023) have measured local variations in the barred-centers of NGC 3351, NGC 3627, and NGC 4321, which were found to be 4-15 times lower compared to the Galactic value. These three galaxies overlap with our sample that host abnormal regions, however, the current available data are not deep enough to detect 13CO in the regions of interest for NGC 3627 and NGC 4321. As for NGC 3351, the locations of high CO/PAH ratio is near the region with very low , located at where bar-driven inflow feeds the nucleus.
Using maps from J. den Brok et al. (2025, ; in preparation) for four galaxies hosting outliers (NGC 1300, NGC 1433, NGC 2566, and NGC 3627), we did not find significant differences in between our abnormal and null-test regions exceeding 1. Y.-H. Teng et al. (2023) have mapped 13CO()/() 12CO/13CO() in NGC 3351 including the outlier region. Figure 5 shows that outlier regions are associated with lower 13CO()/() and higher 12CO/13CO(). These trends suggest that high CO-to-PAH ratio is not due to higher excitation temperature and is consistent with reduced optical depth that allows more CO emission to escape, thereby boosting both the integrated CO intensity and the 12CO/13CO() ratio.
5 Summary
Under the conditions that dust and PAH grains are well mixed with gas and variations in the strength of the interstellar radiation field are modest, stochastically-heated PAH emission is expected to show correlation with the total amount of gas, often traced with CO in molecular gas-dominated galaxies. In line with this expectation, R. Chown et al. (2025) have found a tight and almost linear correlation between 7.7m PAH emission and CO flux on 100 pc scales, across 70 PHANGS-JWST galaxies.
In this work, we present localized departures from this CO–7.7m PAH relation, highlighting regions where the tight correlation breaks down. We have identified regions where the observed CO flux is more than 10 times brighter than that expected from 7.7m PAH emission. We find that such outliers are common, observed in 20 galaxies out of 70 explored in R. Chown et al. (2025). These regions with abnormally high are mostly located along bar lanes and in the centers of barred galaxies with an exception of NGC 1566 where we find such regions to exist in the concave part of the spiral arm as well. Interestingly, outlier regions do not show associated H emission, indicating that the abnormally high CO-to-PAH ratio is not due to massive stars dissociating PAHs (O. V. Egorov et al., 2023; J. Sutter et al., 2024).
@@ 5 Summary
For the JWST Cycle 1 galaxies, we find that the abnormal regions, identified based on elevated CO-to-7.7m PAH ratios, also consistently exhibit elevated ratios when using other PAH (3.3 and 11.3m) and dust continuum bands (10 and 21m). The fact that the ratios are significantly enhanced across multiple PAHs and dust-continuum bands suggests that the abnormal CO-to-PAH ratios are more likely due to overluminous CO emission, rather than suppressed 7.7m PAH emission.
We also find that some abnormal regions either show bigger or smaller PAH size distributions or hints of PAHs being excited by softer radiation fields compared to their null-test regions. However, these trends are not consistent across the system.
@@ 2. High CO emissivity scenario:
Abnormal regions generally show higher CO velocity dispersions () compared to their null-test regions. Higher velocity dispersion leads to lower optical depth in the CO line, thereby increasing the CO emissivity for a given gas mass. This implies that the abnormal regions contains a similar amount of gas as their neighboring regions (as suggested by the similar PAH flux), but the elevated CO emissivity enhances the CO-to-PAH ratio.
Multi-transition CO and isotopologue data enables more direct constraints on opacity and excitation. While can rise with higher excitation, maps for four galaxies show no differences exceeding 1 between abnormal and null-test regions. In NGC 3351, isotopologue ratios show lower and higher in outlier regions, favoring reduced effective optical depth as the driver of the high CO/PAH instead of higher excitation temperature.
@@
In summary, we have explored two possible scenarios to explain the extreme outliers in CO-PAH relation. While the elevated CO velocity dispersion may suggest changes in CO emissivity as the driver of the enhanced CO-to-PAH ratios, this interpretation requires confirmation through deep observations of CO isotopologues and multi-transitions that can directly constrain variations in optical depth and excitation. Future JWST IFU observations targeting additional PAH bands and shock tracers, along with measurements of H2 lines, will be crucial to distinguish between these scenarios.

### v2

@@ IV.1 Suppressed PAH Emission Scenario
As described in Section II.1, the JWST Cycle 1 data provide coverage across multiple filters, enabling measurements of key PAH emission features (at 3.3, 7.7, and 11.3m) as well as the dust continuum (at 10m and 21m). Here, we investigate whether the outlier regions identified in the CO-to-7.7m PAH ratio persist when other PAH or dust continuum dominated bands are used. We also compare PAH properties between the outlier and null-test regions.
IV.1.1 CO-to-mid-IR ratio using other PAH and dust continuum emission
@@
Figure 2 shows the CO flux relative to that expected from other PAH- and dust continuum–dominated bands in NGC 4303, which is one of the galaxies with the most prominent CO-to-7.7m PAH anomalies. We estimate CO fluxes from 3.3 and 11.3m bands using the coefficients from R. Chown et al. (2025a) ( and ), and from 10 and 21m bands using coefficients derived in this work following the same method ( and ).
@@
The outlier regions identified in CO-to-7.7m PAH ratio (red) consistently show elevated ratios across all PAH and continuum bands, with the largest excess observed at 3.3m and the smallest at 10m. Other galaxies hosting outlier regions also show a similar behavior. The significant enhancement across multiple bands indicates that 7.7m is not particularly suppressed and the anomaly is more likely due to overluminous CO emission rather than a physical mechanism that suppresses multiple PAH bands and dust continuum emission at the same time. It is also interesting to note that shows the closest agreement with , even though m traces mostly PAH emission while m traces mostly dust continuum. Further supporting this picture, Appendix A shows three–color HST composites (F814W, F438W, F555W; J. C. Lee et al., 2022) in which the outlier regions display dust extinction similar to their surroundings (see also H. Faustino Vieira et al. 2025). This suggests that outlier regions exhibit comparable dust content compared to their surroundings, making simultaneous suppression of PAH and dust emission, for example due to shielding, an unlikely explanation for the elevated CO/PAH ratios.
IV.1.2 Difference in PAH properties
@@
PAH intensity ratios are sensitive to molecule properties (e.g., size, asymmetry, composition, and charge state) and to local radiation field (S. Hony et al., 2001; F. Galliano et al., 2008; T. S. Y. Lai et al., 2020; A. Maragkoudakis et al., 2020). In particular, 11.3m/3.3m becomes lower as the spectral energy distribution of the illumination radiation becomes harder and PAH size distributions become smaller. The ratio of 11.3m/7.7m also decreases with harder radiation field and is very sensitive to changes of PAH charge variations (T. S. Y. Lai et al., 2022; J. Chastenet et al., 2023; B. T. Draine et al., 2021; B. S. Hensley & B. T. Draine, 2023b; D. Baron et al., 2024, 2025). Furthermore, the ratio of emission in PAH bands to the 21 m dust continuum is considered as a strong tracer of the fraction of dust mass in the form of PAHs (J. Chastenet et al., 2019; J. Sutter et al., 2024; O. V. Egorov et al., 2023).
@@
In Figure 3, we show median and 1 distribution of PAH band ratios to assess whether PAH properties differ between our outlier and null-test regions across Cycle 1 galaxies. The fraction of dust mass in PAHs, , estimated from , shows no significant difference between outlier and null-test regions. We note that the outlier and null-test regions also do not overlap with massive star-forming sites, where is typically suppressed due to PAH destruction by stellar feedback (J. Chastenet et al., 2023; O. V. Egorov et al., 2023, 2025; J. Sutter et al., 2024).
@@
In the middle panel, we compare . While most galaxies follow a one-to-one relation, outlier regions in NGC 1433 and NGC 4535 show somewhat lower compared to their null-test region, suggesting fragmentation into smaller grains or harder radiation field. In contrast, NGC 3627, shows a higher ratio in its outlier region (see also (D. A. Dale et al., 2025)), possibly indicating PAH growth or coagulation onto dust grains, making 7.7m and 11.3m bands fainter (B. T. Draine et al., 2021; B. S. Hensley & B. T. Draine, 2023b). As most of these outlier regions lie along bar lanes with no signs of massive star formation, the differences in suggest that PAH size distributions may be altered by shocks or coagulation as gas and dust flow inward.
@@
Lastly, we compare PAH ratios. The , often used as a tracer of PAH ionization (e.g., J. D. T. Smith et al., 2007; T. S. Y. Lai et al., 2022), has also been shown to correlate with (D. Baron et al., 2024). This implies that the majority of the variation in ratios is driven by the changing shape of the illuminating spectral energy distribution, where the ratio becomes higher as the radiation from old stellar population dominates (high and high ) compared to the radiation from young massive stars (low and low ). In NGC 1433 and NGC 4321, we find somewhat higher in outlier regions ( dex), which may indicate an older stellar population dominating the radiation and/or more neutral PAHs, and therefore the lack of PAH emission compared to CO. However, in most galaxies, outlier and null-test regions agree within 1.
@@
While some outlier regions show hints of altered PAH sizes or hardness of the radiation field, these trends are not consistent across systems. As the PAH ratios explored here are degenerate with multiple physical effects, additional spectroscopic tracers (e.g., H2 rotational and rovibrational lines, [Fe II] for shocks, and a wider suite of PAH features) will be essential to further investigate whether PAH emission is suppressed in these regions.
IV.2 Elevated CO emissivity scenario
@@
Almost all outlier regions lie along the dust lanes in bars, sites where inflowing gas experiences streaming motions and strong shocks due to intersecting orbital families, causing abrupt changes in velocity (W. W. Roberts et al., 1979; E. Athanassoula, 1992; M. C. Sormani et al., 2015; J. A. Sellwood & A. Wilkinson, 1993). These dynamical processes can elevate the CO velocity dispersion and increase the integrated intensity either through line broadening or by reducing the optical depth, which boosts the CO emissivity (A. D. Bolatto et al., 2013; M. Gong et al., 2020; Y.-H. Teng et al., 2023). In addition, shocks and strong shear can raise the excitation temperature and collision rates, populating higher- levels and further enhancing the CO flux for a given molecular gas mass.
@@
In this Section, we test whether the high CO-to-PAH ratios arise from elevated CO emissivity or enhanced excitation by comparing CO velocity dispersions and CO line ratios between the outlier and null-test regions. If the high CO/PAH ratios are associated with higher velocity dispersions or elevated excitation temperature, it would indicate that the -to-PAH ratio between outlier and null-test regions should remain similar and the CO–to– conversion factor () should be at least 10 times lower in the outlier regions compared to the regions with .
IV.2.1 CO velocity dispersion
@@
Figure 4 shows the CO luminosity-weighted average velocity dispersion () as a function of the median CO-to-PAH ratio () of each outlier region and null-test region, across the 20 galaxies hosting CO/PAH outliers. On average, the outlier regions show two times higher and times higher compared to those in null-test regions666When all pixels are considered and not just the null-test region, the distribution of strongly peaks at 1.. The bottom panel shows as a function of the median CO-to-PAH ratio (). We obtain using the Y.-H. Teng et al. (2024) prescription,
with measured within each region. We also compare the distributions of and , where . As outlier regions are associated with higher velocity dispersions and thus lower , the large offset observed between the outlier and null-test regions in distribution decreases when is used. We also show how high (or how low the ) should be in order to explain these outliers solely by enhanced CO-emissivity, under the assumption of a constant -to-PAH ratio (dotted line) between the outlier and null-test regions. While the outlier regions are associated with higher velocity dispersion, it appears it should be even higher (three times more) to explain these outliers from enhanced emissivity.
The bottom panel of Figure 4 shows CO velocity dispersion () as a function of molecular gas surface density () for three galaxies where the enhancement in CO velocity dispersion is prominent. We also divide the data into logarithmically spaced bins in and then compute median and 1 scatter of in each bin for pixels in outlier null-test regions. Within the range of where the outlier and null-test regions overlap, the outlier regions generally exhibit higher , indicating higher CO emissivity that can lead to abnormally high CO-to-PAH ratios. We note that the measured can be inflated by unresolved velocity gradients in the galaxy centers where large-scale rotational shear is steep. However, as we compare outlier regions to nearby null-test regions within the same dynamical structure at similar galactocentric radii, our analysis is expected to be insensitive to beam smearing driven by the large-scale rotation. Moreover, any remaining broadening associated with non-circular motions (e.g., bar-driven streaming, shocks, and shear) is physically relevant to our interpretation.
In Appendix B, the same figure for all 20 galaxies hosting outlier regions is shown. While in some galaxies the velocity dispersions at a given agree within , almost half of them show differences exceeding this level, particularly NGC 1566, NGC 3351, NGC 4321, NGC 2566, NGC 2997, NGC 4569, NGC 4579, and NGC 6300. We also note that when prescription from (Y.-H. Teng et al., 2024) is used, between outlier and null-test regions overlaps much more significantly than when a constant Galactic is adopted (triangles). This again supports the view where high CO-to-PAH ratios are driven by elevated CO emissivity. The outlier and null-test regions contain comparable amounts of PAH emission for a given amount of . These results would imply that the is more than 10 times lower in the outlier region compared to the typical regions with a normal ratio of , suggesting a strong sub-kpc variation in as observed in the Milky Way and centers of barred galaxies (L. Blitz et al., 1985; Y.-H. Teng et al., 2022, 2023).
IV.2.2 CO excitation from multi-transition lines and CO isotopologues
@@
Using CO isotopologues observations in their low rotational transitions, Y.-H. Teng et al. (2022) and Y.-H. Teng et al. (2023) have measured local variations in the barred-centers of NGC 3351, NGC 3627, and NGC 4321, which were found to be 4-15 times lower compared to the Galactic value. These three galaxies overlap with our sample that host outlier regions, however, the current available data are not deep enough to detect 13CO in the regions of interest for NGC 3627 and NGC 4321. As for NGC 3351, the locations of high CO/PAH ratio is near the region with very low , located at where bar-driven inflow feeds the nucleus.
Using maps from J. den Brok et al. (2025, J. den Brok et al. in preparation) for four galaxies hosting outliers (NGC 1300, NGC 1433, NGC 2566, and NGC 3627), we did not find significant differences in between our outlier and null-test regions exceeding 1. Y.-H. Teng et al. (2023) have mapped 13CO()/() 12CO/13CO() in NGC 3351 including the outlier region. Figure 5 shows that outlier regions are associated with lower 13CO()/() and higher 12CO/13CO(). These trends suggest that the high CO-to-PAH ratio is not due to higher excitation temperature and is consistent with reduced optical depth that allows more CO emission to escape, thereby boosting both the integrated CO intensity and the 12CO/13CO() ratio.
V Summary
Under the conditions that dust and PAH grains are well mixed with gas and variations in the strength of the interstellar radiation field are modest, stochastically-heated PAH emission is expected to show correlation with the total amount of gas, often traced with CO in molecular gas-dominated galaxies. In line with this expectation, R. Chown et al. (2025a) have found a tight and almost linear correlation between 7.7m PAH emission and CO flux on 100 pc scales, across 70 PHANGS-JWST galaxies.
In this work, we present localized departures from this CO–7.7m PAH relation, highlighting regions where the tight correlation breaks down. We have identified regions where the observed CO flux is more than 10 times brighter than that expected from 7.7m PAH emission. We find that such outliers are common, observed in 20 galaxies out of 70 explored in R. Chown et al. (2025a). These regions with abnormally high are mostly located along bar lanes and in the centers of barred galaxies with an exception of NGC 1566 where we find such regions to exist in the concave part of the spiral arm as well. Interestingly, outlier regions do not show associated H emission, indicating that the high CO-to-PAH ratio is not due to massive stars dissociating PAHs (O. V. Egorov et al., 2023; J. Sutter et al., 2024).
@@ V Summary
For the JWST Cycle 1 galaxies, we find that the outlier regions, identified based on elevated CO-to-7.7m PAH ratios, also consistently exhibit elevated ratios when using other PAH (3.3 and 11.3m) and dust continuum bands (10 and 21m). The fact that the ratios are significantly enhanced across multiple PAHs and dust-continuum bands suggests that the high CO-to-PAH ratios are more likely due to overluminous CO emission, rather than suppressed 7.7m PAH emission.
We also find that some outlier regions either show bigger or smaller PAH size distributions or hints of PAHs being excited by softer radiation fields compared to their null-test regions. However, these trends are not consistent across the system.
@@ 2. High CO emissivity scenario:
Outlier regions generally show higher CO velocity dispersions () compared to their null-test regions. Higher velocity dispersion leads to lower optical depth in the CO line, thereby increasing the CO emissivity for a given gas mass. This implies that the outlier regions contains a similar amount of gas as their neighboring regions (as suggested by the similar PAH flux), but the elevated CO emissivity enhances the CO-to-PAH ratio.
Multi-transition CO and isotopologue data enables more direct constraints on opacity and excitation. While can rise with higher excitation, maps for four galaxies show no differences exceeding 1 between outlier and null-test regions. In NGC 3351, isotopologue ratios show lower and higher in outlier regions, favoring reduced effective optical depth as the driver of the high CO/PAH instead of higher excitation temperature.
@@
In summary, we have explored two possible scenarios to explain the extreme outliers in CO-PAH relation. While the elevated CO velocity dispersion may suggest changes in CO emissivity as the driver of the enhanced CO-to-PAH ratios, this interpretation requires confirmation through deep observations of CO isotopologues and multi-transitions that can directly constrain variations in optical depth and excitation. Future JWST IFU observations targeting additional PAH bands, [FeII] shock tracers, Hydrogen recombination lines, and H2 rotational and rovibrational lines, will help constrain the physical conditions of the PAHs, identify shocks which can destroy PAHs, and map the warm molecular gas, thereby distinguishing whether the CO emission is enhanced by the local environment.